OpenStudy (anonymous):

Find the LCM z^3+8z^2+16z, z^2-8z

OpenStudy (anonymous):

z^3+8z^2+16z = z(z^2+8z+16)=z(z+4)(z+4) z^2-8z = z(z-8) So LCM is z(z-8)(z+4)^2

OpenStudy (anonymous):

thank you!